A physics student does 0.0273 J of work on a cart to push it 5.80 cm along on air track. If the acceleration of the cart was 0.314 m/s2, what was the mass of the car? 

 

  

 

 

 

 



1
Expert's answer
2021-06-02T09:37:06-0400

Let's first find the force that the student used to push the cart along on air track:


"W=Fd,""F=\dfrac{W}{d}=\dfrac{0.0273\ J}{0.058\ m}=0.47\ N."

Finally, we can find the mass of the car from the Newton's Second Law of Motion:


"F=ma,""m=\dfrac{F}{a}=\dfrac{0.47\ N}{0.314\ \dfrac{m}{s^2}}=1.5\ kg."